OpenStudy (anonymous):

i need someone to walk me through solving this equation: 200x + 1,000 = 10(1.5^x) medal and fan.(:

OpenStudy (anonymous):

Is that 1.5 to the x power?

OpenStudy (anonymous):

yes

OpenStudy (sweetburger):

lol i plugged this into my calculator and it took about a min to solve for an x value

OpenStudy (anonymous):

haha

OpenStudy (sweetburger):

well first i know we might aswell divide everything by 10 to get rid of it

Nnesha (nnesha):

first of all divide both sides by 10 so you can take ln both sides.......-....

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

one second please

OpenStudy (anonymous):

okay i got it

OpenStudy (anonymous):

sorry it took so long, i had to deal with something else.

OpenStudy (sweetburger):

you can factor out a 20 on the right side of the equation

OpenStudy (anonymous):

i have 200x + 100 = 1(1.5^x)

OpenStudy (anonymous):

why am i putting a twenty there?

OpenStudy (sweetburger):

ok so u didnt factor out the 10 from everything

OpenStudy (anonymous):

yes i divided 10 and 1000 by 10

OpenStudy (sweetburger):

should be 20x + 100 = 1.5^x

OpenStudy (anonymous):

wait what

OpenStudy (sweetburger):

divide everything by 10...

OpenStudy (anonymous):

i thought you only divide the sides that don't have x

OpenStudy (anonymous):

so 1000 and 10

OpenStudy (sweetburger):

what... u have confused me

OpenStudy (sweetburger):

just divide each value in the equation by 10 so 200/10 1000/10 and 10/10

OpenStudy (sweetburger):

200x/10 +1000/10 = 10/10(1.5^x) now solve that @grunge101

OpenStudy (anonymous):

i thought you weren't supposed to divide the numbers with x

OpenStudy (sweetburger):

why did u think that? remember that 200 in front of x is just a coefficient you still have to reduce it

OpenStudy (anonymous):